### v1

Abstract
We present the discovery of 11 new transiting brown dwarfs and low-mass M-dwarfs from NASA’s TESS mission: TOI-2844, TOI-3122, TOI-3577, TOI-3755, TOI-4462, TOI-4635, TOI-4737, TOI-4759, TOI-5240, TOI-5467, and TOI-5882. They consist of 5 brown dwarf companions and 6 very low mass stellar companions ranging in mass from 25 to 128 . We used a combination of photometric time-series, spectroscopic, and high resolution imaging follow-up as a part of the TESS Follow-up Observing Program (TFOP) in order to characterize each system. With over 50 transiting brown dwarfs confirmed, we now have a large enough sample to directly test different formation and evolutionary scenarios. We provide a renewed perspective on the transiting brown dwarf desert and its role in differentiating between planetary and stellar formation mechanisms. Our analysis of the eccentricity distribution for the transiting brown dwarf sample does not support previous claims of a transition between planetary and stellar formation at 42 . We also contribute a first look into the metallicity distribution of transiting companions in the range , showing that this too does not support a 42 transition. Finally, we also detect a significant lithium absorption feature in one of the brown dwarf hosts (TOI-5882) but determine that the host star is likely old based on rotation, kinematic, and photometric measurements. We therefore claim that TOI-5882 may be a candidate for planetary engulfment.
1 Introduction
Since the launch of NASA’s Transiting Exoplanet Survey Satellite (TESS) in 2018 (Ricker et al. 2015), the number of brown dwarfs (BDs) known to transit their host stars has increased rapidly from just 16 systems to . These BDs, which are defined as objects within the mass range of 13 - 80 , fuse only deuterium in their cores. This differentiates them from planets, which undergo no fusion, and stars, which ignite hydrogen fusion. However, the deuterium and hydrogen burning limits have been shown to be less clear than this definition would imply. Spiegel et al. (2011) showed that the lower limit varies from 11 - 16 , while Baraffe et al. (2003) showed that hydrogen fusion can ignite between 75 - 80 . The spread in both of these estimates can be explained by variation in the chemical composition and formation conditions of the BD. While these definitions provide insight into the physical processes taking place in BD interiors, they offer little insight into how they form.
Reframing our perspective on BDs into one motivated by formation and evolution has long been advocated for by some members of the BD community (Chabrier et al. 2014; Burrows 2014; Carmichael et al. 2021), where objects would be distinguished based on whether they form through a planet-like or a star-like formation mechanism. BDs forming like planets would undergo a core accretion pathway (Pollack et al. 1996), commonly referred to as a ”bottom-up” approach. The star-like BDs on the other hand would form via direct gravitational collapse, or ”top-down”, which can happen either within the circumstellar disk or at the core scale (Adams et al. 1989; Bate 2012; Kratter & Lodato 2016). Differentiating between these two formation pathways remains challenging, since it is unclear under what conditions each mechanism dominates, and whether there are any observable parameters that could distinguish them. Fortunately, in the era of TESS, we have begun to accumulate transiting BDs en masse, allowing us to pursue the question of BD formation from a different perspective. This budding population of transiting BDs is particularly enticing for studying BD formation because it provides a complementary, and in many cases, more complete understanding of the BD compared to the previously studied objects, which have primarily been discovered via direct imaging or radial velocity (RV) techniques. The transiting population serves as a complementary dataset to these other populations because transits provide a direct measurement of BD radii, a property which often can only be otherwise inferred with evolutionary models based on the observed spectrum and luminosity. This measurement is vital because BDs tend to contract with age, while also decreasing in size as mass increases (Baraffe et al. 2003; Saumon & Marley 2008; Burrows et al. 2001; Phillips et al. 2020). Thus there exists a degeneracy between mass, radius, and age for BDs making it difficult to test the substellar models with observed systems unless all three variables can be measured. These transiting systems provide direct, independent measurements on two of these degenerate parameters, and in cases where the host star’s age can be precisely determined, all three (e.g. Gillen et al. 2017; Nowak et al. 2017; David et al. 2019; Vowell et al. 2023).
This rapidly growing population of transiting BDs also allows us to revisit the longstanding idea of the so-called ”brown dwarf desert”. Prior work has shown a dearth of brown dwarfs orbiting main sequence host stars with semi-major axes au. (Marcy et al. 1997; Latham et al. 1998). Ma & Ge (2014) refined our understanding of the brown dwarf desert by investigating the population of all published brown dwarfs discovered with the RV method at the time. Here they found that the ”driest land” of the desert lies between and with period days). The authors attribute this feature to being a result of different formation mechanisms dominating in different mass regimes. Namely, that stellar binary formation is responsible for the systems with BD companions while formation in the protoplanetary disk explains the systems with BDs . However, the sample in this study with period days was quite small at only 25 brown dwarfs. Furthermore, by virtue of being a RV study, it was restricted to only probing rather than the BD mass directly, unable to break the degeneracy in most cases, a complication that the transiting population does not have.
As this population of transiting brown dwarfs expanded in the era of space-based transit surveys, several new discoveries noted an ”oasis” forming in the desert (Carmichael et al. 2020; Šubjak et al. 2020; Henderson et al. 2024a) with new transiting systems beginning to populate the driest region of the desert noted by Ma & Ge (2014). With the new discoveries presented in this work, the transiting BD population now exceeds 50 systems, more than double the size of the population Ma & Ge (2014) had access to, opening the door for a reevaluation of brown dwarf desert from a new perspective. Hence, in this paper we present the discovery of 11 new transiting companions from NASA’s TESS mission. 6 of these systems are BD, with 3 lying within the Ma & Ge (2014) defined ”driest” region of the BD desert. We confirmed the remaining 6 non-BD, companions as very low-mass stars . In §2 of this manuscript, we present all the observations collected for each system in this work. §3 details our analysis of each system using EXOFASTv2 (Eastman et al. 2013, 2019). In §4 we provide a discussion on how these new systems fit into the population as a whole with a renewed perspective on the BD desert. We also discuss a detection of lithium in the host star of the BD companions presented here (TOI-5882). Finally, we present our conclusions in §5.
2 Observations
In the following subsections we present all observations collected and analyzed for each target in this sample. To briefly summarize, each target has a suite of observations that serve to characterize the host star and/or companion and rule out false-positive scenarios. Generally, these observations include archival multiband observations from various ground-based missions, time-series photometry from both space and ground-based telescopes, spectroscopy, and high resolution imaging. See Table 1 for the relevant results of the archival data associated with each system.
TOI-2844 | TOI-3122 | TOI-3577 | TOI-3755 | Source | ||
Other identifiers: | ||||||
TESS Input Catalog | TIC 387342052 | TIC 61117473 | TIC 396133015 | TIC 281196902 | ||
TYCHO-2 | TYC 771-367-1 | TYC 6773-1-1 | TYC 3608-647-1 | — | ||
2MASS | J07204878+1301073 | J15074899-2809237 | J21482300+4820042 | J04385936+6640161 | ||
Gaia DR3 | 3166196736096450816 | 6212565847439064192 | 1977894600881987328 | 483359575160953728 | ||
Astrometric Parameters: | ||||||
Right Ascension (h:m:s) | 07:20:48.778 | 15:07:48.989 | 21:48:23.012 | 04:38:59.373 | 1 | |
Declination (d:m:s) | 13:01:07.359 | -28:09:23.782 | 48:20:04.355 | 66:40:16.177 | 1 | |
Gaia DR3 proper motion in RA (mas yr-1) | 1 | |||||
Gaia DR3 proper motion in Dec (mas yr-1) | 1 | |||||
Gaia DR3 Parallax (mas) | 1 | |||||
Projected rotational velocity (km s-1) | 2 | |||||
Photometric Parameters: | ||||||
Gaia mag. | 1 | |||||
Gaia mag. | 1 | |||||
Gaia mag. | 1 | |||||
TESS mag. | 3 | |||||
2MASS mag. | 4 | |||||
2MASS mag. | 4 | |||||
2MASS mag. | 4 | |||||
WISE mag. | 5 | |||||
WISE mag. | 5 | |||||
WISE mag. | 5 |
TOI-4462 | TOI-4635 | TOI-4737 | TOI-4759 | Source | ||
Other identifiers: | ||||||
TESS Input Catalog | TIC 76420654 | TIC 337129672 | TIC 142532090 | TIC 49705089 | ||
TYCHO-2 | TYC 2635-1030-1 | — | — | — | ||
2MASS | J18184078+3615175 | J02143112+0804481 | J06533851-1326106 | J06234422-2401288 | ||
Gaia DR3 | 4605954852723545088 | 2521579495665163008 | 2949605211853441664 | 2936390357694302336 | ||
Astrometric Parameters: | ||||||
Right Ascension (h:m:s) | 18:18:40.777 | 02:14:31.26 | 06:53:38.51 | 06:23:44.23 | 1 | |
Declination (d:m:s) | 36:15:17.53 | 08:04:45.34 | -13:26:10.662 | -24:01:28.877 | 1 | |
Gaia DR3 proper motion in RA (mas yr-1) | 1 | |||||
Gaia DR3 proper motion in Dec (mas yr-1) | 1 | |||||
Gaia DR3 Parallax (mas) | 1 | |||||
Projected rotational velocity (km s-1) | 2 | |||||
Photometric Parameters: | ||||||
Gaia mag. | 1 | |||||
Gaia mag. | 1 | |||||
Gaia mag. | 1 | |||||
TESS mag. | 3 | |||||
2MASS mag. | 4 | |||||
2MASS mag. | 4 | |||||
2MASS mag. | 4 | |||||
WISE mag. | 5 | |||||
WISE mag. | 5 | |||||
WISE mag. | 5 |
TOI-5240 | TOI-5467 | TOI-5882 | Source | ||
Other identifiers: | |||||
TESS Input Catalog | TIC 40055053 | TIC 83275782 | TIC 232941965 | ||
TYCHO-2 | TYC 2663-268-1 | — | TYC 2695-1754-1 | ||
2MASS | J19322010+3456254 | J06173449+2826431 | J20473329+3444151 | ||
Gaia DR3 | 2046792606517797632 | 3433414139371114368 | 1869489729418662528 | ||
Astrometric Parameters: | |||||
Right Ascension (h:m:s) | 19:32:20.11 | 06:17:34.49 | 20:47:33.291 | 1 | |
Declination (d:m:s) | 34:56:25.423 | 28:26:43.117 | 34:44:15.238 | 1 | |
Gaia DR3 proper motion in RA (mas yr-1) | 1 | ||||
Gaia DR3 proper motion in Dec (mas yr-1) | 1 | ||||
Gaia DR3 Parallax (mas) | 1 | ||||
Projected rotational velocity (km s-1) | 2 | ||||
Photometric Parameters: | |||||
Gaia mag. | 1 | ||||
Gaia mag. | 1 | ||||
Gaia mag. | 1 | ||||
TESS mag. | 3 | ||||
2MASS mag. | 4 | ||||
2MASS mag. | 4 | ||||
2MASS mag. | 4 | ||||
WISE mag. | 5 | ||||
WISE mag. | 5 | ||||
WISE mag. | 5 |
NOTES:
The uncertainties of the photometric measurements have a systematic floor applied that is usually larger than the reported catalog errors.
Right Ascension and Declination are in epoch J2000. The coordinates come from Vizier where the Gaia RA and Dec have been precessed and corrected to J2000 from epoch J2016.
Sources: (1) Gaia Collaboration et al. (2023); (2) §2.3; (3) Stassun et al. (2019); (4) Cutri et al. (2003); Skrutskie et al. (2006); (5) Wright et al. (2010); Cutri et al. (2012)
2.1 TESS Photometry
Each system presented here initially showed signs of an orbiting companion via transits detected by TESS. TESS has a mosaic of four CCD cameras each with a x field of view, and a pixel size of 21”. In combination, this makes the TESS field of view x for each sector, which is observed for approximately 27 days before moving to a new sector of sky. TESS observes at a 2-second, and in the TESS prime mission, the data were processed into 2-minute stacks for select stars, with the rest of the field being processed at 30-minute cadence. This prime mission observed 80% of the entire sky with the largest gaps in coverage occurring near the ecliptic plane. As TESS transitioned to its first, and now second, extended missions it continues to observe even more of the ecliptic plane. In this second extended mission, most preselected targets are now processed at 120-seconds while a smaller number are processed at 20-second cadence. Full-frame images are processed at 200-seconds.
The systems presented here were observed by TESS between Sectors with cadences ranging from 30-minutes in the prime mission to as low as 2-minutes in the extended mission. The TESS data were originally downloaded and reduced using both the TESS Science Processing Operations Center (SPOC) Pipeline (Jenkins et al. 2016) and the MIT Quick-Look Pipeline (QLP; Huang et al. 2020a, b; Kunimoto et al. 2021). The initial detection of a transit-like signal was discovered and vetted by the faint-star QLP search (Kunimoto et al. 2022) for 10 out the 12 systems presented in this paper, all but TOI-4462 and TOI-4635. These two TOIs were initially detected by the QLP and SPOC pipelines respectively, and then vetted by the TESS Science Office. The diagnostic tests described in Twicken et al. (2018) were used to evaluate whether the transit-like signal is indeed Keplerian. Upon passing, each system was designated as a TESS Object of Interest (TOI; Guerrero et al. 2021). While both the QLP and SPOC pipelines correct for contamination by known nearby stars, we choose to use the SPOC light curves with the shortest cadence in our analysis wherever possible for consistency. See Table 2 for full details on the sectors, cadence, and pipeline used for each source. It should be noted that all QLP light curves are processed from the full frame images. The TESS-SPOC (Caldwell et al. 2020) light curves are produced by the SPOC on a best-effort basis under the leadership of Doug Caldwell, the PI, and delivered as high-level science products to the Mikulski Archive for Space Telescopes (MAST) rather than as part of the official mission data products produced by the SPOC.
We downloaded the individual light curves from the MAST using the lightkurve111https://github.com/lightkurve/lightkurve package (Lightkurve Collaboration et al. 2018). We then removed any long-term variability (both stellar and instumental) by fitting a spline to the flux and dividing the light curve by the best fitting spline model. We used the Keplerspline222https://github.com/avanderburg/keplerspline package for this process as described in (Vanderburg & Johnson 2014). We remove most of the out-of-transit data, electing to keep just one transit duration of baseline on either side of the transit.
Target | TESS Sector | Cadence (s) | Pipeline |
TOI-2844 | 7 | 1800 | TESS-SPOC |
— | 33 | 600 | TESS-SPOC |
— | 44 | 600 | TESS-SPOC |
— | 45 | 600 | TESS-SPOC |
— | 46 | 600 | TESS-SPOC |
— | 71 | 120 | SPOC |
— | 72 | 120 | SPOC |
TOI-3122 | 11 | 1800 | QLP |
— | 38 | 600 | QLP |
— | 65 | 120 | SPOC |
TOI-3577 | 8 | 1800 | QLP |
— | 56 | 120 | SPOC |
— | 76 | 120 | SPOC |
TOI-3755 | 19 | 1800 | TESS-SPOC |
— | 59 | 120 | SPOC |
— | 73 | 120 | SPOC |
TOI-4462 | 26 | 1800 | TESS-SPOC |
— | 40 | 600 | TESS-SPOC |
— | 53 | 600 | TESS-SPOC |
— | 54 | 600 | TESS-SPOC |
— | 74 | 120 | SPOC |
TOI-4635 | 42 | 120 | SPOC |
— | 43 | 120 | SPOC |
— | 70 | 120 | SPOC |
— | 71 | 120 | SPOC |
TOI-4737 | 6 | 1800 | TESS-SPOC |
— | 7 | 1800 | QLP |
— | 33 | 600 | TESS-SPOC |
TOI-4759 | 6 | 1800 | QLP |
— | 33 | 600 | QLP |
TOI-5240 | 14 | 1800 | QLP |
— | 40 | 600 | QLP |
— | 41 | 600 | TESS-SPOC |
— | 54 | 600 | QLP |
— | 55 | 600 | QLP |
— | 74 | 120 | SPOC |
— | 75 | 120 | SPOC |
TOI-5467 | 43 | 600 | TESS-SPOC |
— | 44 | 600 | TESS-SPOC |
— | 45 | 600 | TESS-SPOC |
— | 71 | 120 | TESS-SPOC |
— | 72 | 120 | TESS-SPOC |
TOI-5882 | 15 | 1800 | TESS-SPOC |
— | 41 | 600 | TESS-SPOC |
— | 55 | 600 | TESS-SPOC |
— | 75 | 120 | SPOC |
2.2 Ground-Based Time-Series Photometry
In order to confirm that the signal observed by TESS is on-target not originating from a nearby eclipsing binary that is blended with the target star, we gathered ground-based, time-series photometry of each system as the companion transited its host star. Since TESS has a relatively large pixel scale (21″ per pixel), the shallow eclipses we measured, which are consistent with roughly 1 , can be easily mimicked when a different, nearby eclipsing binary happens to fall on the same photometric aperture as the target star. The much deeper eclipses of the nearby eclipsing binary become diluted by the target star to mimic a much shallower event. Seeing-limited ground-based telescopes can have a much higher angular resolution than TESS typically 1-2″ and therefore can confirm that the signal is on-target, thereby ruling out nearby eclipsing binaries at all but the closest separations. It also has the benefit of observing in multiple wavelengths to confirm that the transit-like signal is achromatic. This is helpful because the eclipse depth of an eclipsing binary is nearly always wavelength dependent since the occulting body cannot be treated as a non-luminous sphere. We also note here, that while 6 of the companions presented here are low-mass M-dwarfs, and hence are eclipsing binaries themselves, they are so low in mass that we can still treat them as black spheres since they contribute negligibly to the overall flux of the system (Stevens et al. 2018). This process no only rules out the nearby eclipsing binary false positive, but also serves to refine the ephemerides of systems in which these data are able to extend the photometric baseline.
The observations for these systems were collected through the TESS Follow-up Observing Program (TFOP; Collins et al. 2018) from various observatories as shown in Table 3. The Las Cumbres Observatory Global Telescope (LCOGT; Brown et al. 2013) was responsible for 14 light curves from the following sites: McDonald Observatory (McD), Teide Observatory (TEID), South African Astronomical Observatory (SAAO), and Cerro Tololo Inter-American Observatory (CTIO). The remaining light curves were contributed by the following facilities: Calar Alto Observatory, Brierfield Observatory, the Telescopio Carlos Sánchez (TCS) at Teide Observatory, Grand-Pra (GdP) Observatory, Thacher Observatory (Swift et al. 2022), KeplerCam at the Fred Lawrence Whipple Observatory (FLWO), and the Acton Sky Portal.
All data sets, except for the observations of TOI-3577 from MuSCAT2, were reduced and their light curves extracted using AstroImageJ (AIJ; Collins et al. 2017). To do this, we use AIJ’s multi-aperture photometry tool using at least five similarly bright comparison stars. We use AIJ’s built-in transit fitting tool to assess the quality of the data and determine detrending parameters. Generally, we only detrend against the parameters that strongly correlate with the apparent brightness of the copmanion stars as they change over the course of the night. We also take care not to be too aggressive in our detrending, choosing to adopt detrending only when the Bayesian Information Criterion of AIJ’s transit-only fit significantly favors the detrended model. The detrending parameters used for each light curve can be found in Table 3. See §D in the appendix of Collins et al. (2017) for a detailed description of each detrending parameter. Finally, we normalized the data to the out-of-transit baseline and incorporated each light curve (with detrending) into our global fitting process (see §3).
Our follow-up observations of TOI-3577 were taken by MuSCAT2 on the TCS from Teide Observatory in Tenerife, Spain (Narita et al. 2019). MuSCAT2 is a multi-band imager with four cameras, each with a field of view of . This set-up allows for simultaneous observation in multiple bands, which in our case, were the , , , and bands. These data were reduced by the dedicated MuSCAT2 pipeline (Parviainen et al. 2019), and incorporated into our global fit.
=-0.9in TIC ID TOI Telescope Camera Observation Date (UT) Telescope Size (m) Filter Pixel Scale (arcsec) Exposure Time (s) Detrend params 387342052 2844 LCO-McD QHY600 2023 April 5 0.35 i’ 0.7 135 Airmass Zeiss Calar Alto iKon-XL 230 2023 November 21 1.23 R 0.314 90 None LCO-TEID Sinistro 2023 November 29 1.0 i’ 0.389 19 None 61117473 3122 Brierfield Moravian 16803 2023 May 15 0.36 R 0.735 180 Airmass 396133015 3577 TCS MuSCAT2 2023 July 16 1.52 g’ 0.44 10 None TCS MuSCAT2 2023 July 16 1.52 r’ 0.44 5 None TCS MuSCAT2 2023 July 16 1.52 i’ 0.44 5 None TCS MuSCAT2 2023 July 16 1.52 z_s 0.44 10 None 281196902 3755 GdP FLI4710 2022 March 3 0.4 i’ 0.73 90 None Thacher CDK-700 Teledyne PIXIS 2023 October 18 0.7 r’ 0.608 40 None 76420654 4462 FLWO KeplerCam 2024 March 19 1.2 i’ 0.672 6 Airmass LCO-TEID Sinistro 2024 April 3 1.0 i’ 0.389 38 tot_C_cnts 337129672 4635 LCO-SAAO Sinistro 2023 November 22 1.0 z’ 0.389 37 None LCO-CTIO-fa04 Sinistro 2023 December 5 1.0 z’ 0.389 37 None LCO-CTIO-fa15 Sinistro 2023 December 5 1.0 z’ 0.389 37 Airmass 142532090 4737 LCO-TEID Sinistro 2023 November 26 1.0 i’ 0.389 33 None LCO-CTIO Sinistro 2023 December 6 1.0 i’ 0.389 33 None LCO-SAAO Sinistro 2023 December 24 1.0 i’ 0.389 33 Airmass 49705089 4759 LCO-SAAO-fa06 Sinistro 2024 February 5 1.0 i’ 0.389 44 Airmass LCO-SAAO-fa14 Sinistro 2024 February 5 1.0 i’ 0.389 44 None 40055053 5240 LCO-TEID Sinistro 2023 August 2 1.0 i’ 0.389 26 None 83275782 5467 FLWO KeplerCam 2023 March 4 1.2 i’ 0.672 15 Airmass Acton Sky Portal SBIG A4710 2023 March 20 0.36 r’ 1.0 20 Airmass LCO-McD Sinistro 2023 October 13 1.0 i’ 0.389 29 tot_C_cnts 232941965 5882 LCO-McD Sinistro 2023 June 16 1.0 z’ 0.389 45 None
NOTE: All lightcurves are available on ExoFOP333https://exofop.ipac.caltech.edu.
2.3 TRES Spectroscopy
We collected spectroscopic observations for each system to measure the mass and eccentricity of their companions while also further ruling out the false-positive scenario of nearby eclipsing binaries. While several of the systems presented here have companions above the hydrogen-burning boundary, and thus are eclipsing binaries themselves, none of them have companions that are bright enough to be detected photometrically or spectroscopically. Hence, they are all single-lined spectroscopic binaries. Any potential nearby eclipsing binaries (both bound and unbound) that cannot be ruled out by ground-based photometry can be ruled out by spectroscopy within the angular diameter of the fiber. These are ruled out by the fact that the companions presented here are significantly more massive than their giant planet counterparts. The Doppler motion of the host stars’ spectral lines is too large to be mimicked by a nearby eclipsing binary without resolving a secondary set of spectral features.
We obtained the spectroscopic measurements for each system in this sample via the Tillinghast Reflector Échelle Spectrograph (TRES) on the 1.5-meter Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory on Mt. Hopkins, Arizona. The TRES instrument is a fiber-fed, échelle spectrograph with a resolving power of 44,000. We reduced the spectra according to Buchhave et al. (2010) and analyzed each observation with the Stellar Parameter Classification (SPC) tool (Buchhave et al. 2012) in order to measure the metallicity, effective temperature, surface gravity, and projected rotational velocity of the star. We incorporated the average metallicity for each system into our analysis as a Gaussian prior in our global fits (see §3). We did not incorporate the effective temperature or surface gravity measurements from SPC as priors because these quantities are better constrained by the fit itself. This is due to the fact that EXOFASTv2 simultaneously models the spectral energy distribution, companion’s transit, and stellar evolutionary models (Eastman et al. 2023)
Finally, we derived the radial velocities according to the methods described in Quinn et al. (2012), except that we do not cross-correlate against a template spectrum. Instead, we create a high S/N, median-combined observed spectrum that we cross-correlate with each individual spectrum. See Table 4 for a sample radial velocity point for each system (the full table of radial velocities is available in machine-readable form in the online journal).
Target | RV (m s-1) | (m s-1) | |
TOI-2844 | 2459528.8676 | -6696 | 1791 |
TOI-3122 | 2459651.9406 | 14841 | 145 |
TOI-3577 | 2459395.9184 | 52 | 70 |
TOI-3755 | 2459477.9662 | -8617 | 31 |
TOI-4462 | 2459468.6623 | 944 | 76 |
TOI-4635 | 2459556.7367 | 6170 | 19 |
TOI-4737 | 2459583.9325 | 2529 | 65 |
TOI-4759 | 2459623.7415 | -64 | 53 |
TOI-5240 | 2459681.9612 | 20204 | 375 |
TOI-5467 | 2459697.6358 | -237 | 352 |
TOI-5882 | 2459899.6240 | 108 | 55 |
NOTE: The full table of RVs for each system is available in machine-readable form in the online journal.
2.4 High Resolution Imaging
Target | Telescope | Instrument | Image Type | Filter | Contrast | Observation Date (UT) | Detection?† |
TOI-2844 | SOAR (4.1 m) | HRCam | Speckle | 5.8 mag at 1 | 2022 Apr 15 | No | |
— | WIYN (3.5 m) | NESSI | Speckle | 832 nm | — | 2022 Apr 18 | No |
TOI-3122 | SOAR (4.1 m) | HRCam | Speckle | 5.0 mag at 1 | 2022 Apr 25 | No | |
TOI-3577 | Palomar (5.0 m) | PHARO | AO | Br | 5.8 mag at 0.5 | 2023 Jun 7 | No |
TOI-3755 | SAI (2.5 m) | Speckle Polarimeter | Speckle | 5.7 mag at 1 | 2021 Oct 29 | No | |
TOI-4462 | SAI (2.5 m) | Speckle Polarimeter | Speckle | 6.3 mag at 1 | 2023 Jan 22 | Yes | |
— | SAI (2.5 m) | Speckle Polarimeter | Speckle | 5.3 mag at 1 | 2024 Feb 24 | Yes | |
— | Palomar (5.0 m) | PHARO | AO | 7.2 mag at 0.5 | 2024 Apr 21 | Yes | |
— | Palomar (5.0 m) | PHARO | AO | 6.9 mag at 0.5 | 2024 Apr 21 | Yes | |
TOI-4635 | Shane (3.0 m) | ShARCS | AO | — | 2021 Nov 21 | No | |
— | Shane (3.0 m) | ShARCS | AO | — | 2021 Nov 21 | No | |
— | SOAR (4.1 m) | HRCam | Speckle | 6.7 mag at 1 | 2024 Jan 8 | No | |
TOI-4737 | Gemini (8.0 m) | Zorro | Speckle | nm | 4.3 mag at 0.5 | 2022 Mar 19 | No |
— | Gemini (8.0 m) | Zorro | Speckle | nm | 6.0 mag at 0.5 | 2022 Mar 19 | No |
— | SOAR (4.1 m) | HRCam | Speckle | 5.6 mag at 1 | 2022 Apr 15 | No | |
TOI-4759 | SOAR (4.1 m) | HRCam | Speckle | 6.2 mag at 1 | 2022 Apr 15 | No | |
TOI-5240 | Palomar (5.0 m) | PHARO | AO | Br | 6.7 mag at 0.5 | 2023 Jun 6 | Yes |
— | SAI (2.5 m) | Speckle Polarimeter | Speckle | 6.3 mag at 1 | 2023 Sep 1 | No | |
TOI-5467 | SAI (2.5 m) | Speckle Polarimeter | Speckle | 6.2 mag at 1 | 2022 Dec 12 | No | |
TOI-5882 | Palomar (5.0 m) | PHARO | AO | Br | 6.8 mag at 0.5 | 2023 Jun 6 | No |
— | SAI (2.5 m) | Speckle Polarimeter | Speckle | 7.4 mag at 1 | 2023 Aug 28 | No |
NOTE: All images and contrast curves are available on ExoFOP.
Detection refers to a positive detection of a star within the field of view of the AO or speckle instrument, subject to the maximum contrast possible with the instrument in question.
While ground-based transits rule out nearby eclipsing binaries at most scales, if another source is close enough to the target star, it may be blended both in TESS and from the ground. Therefore, to verify that there is no contamination at these very small separations, and in order to detect any potentially bright companions, we utilized high-resolution imaging. We employed both Adaptive Optics (AO) and speckle imaging instruments to obtain our high resolution images for these systems.
We used the ShARCS and PHARO instruments for AO imaging. The ShARCS instrument is on the Shane 3.0-meter telescope located at Lick Observatory (Kupke et al. 2012; Gavel et al. 2014; McGurk et al. 2014). The PHARO instrument is on the Palomar Hale 5-meter telescope at Palomar Observatory (Hayward et al. 2001). For our speckle observations we used the following telescopes and instruments: HRCam on the Southern Astrophysical Research (SOAR) 4.1-meter telescope at CTIO (Tokovinin 2018; Ziegler et al. 2020), the NN-EXPLORE Exoplanet Stellar Speckle Imager (NESSI; Scott et al. 2018) on the WIYN 3.5-meter telescope at Kitt Peak Observatory, the Speckle Polarimeter on the 2.5-meter telescope at the Caucasian Mountain Observatory of the Sternberg Astronomical Institute (SAI) at Lomonosov Moscow State University, and the Zorro instrument on the Gemini-South 8-meter telescope. The Speckle Polarimeter used an Andor iXon 897 Electron Multiplying CCD for the observation of TOI-3755 (Safonov et al. 2017). All other observations from this instrument used a Hamamatsu ORCA-quest CMOS detector (Strakhov et al. 2023). See Table 5 for a summary of each observation including the dates each system was observed, filters used, contrast achieved, and whether a nearby companion was detected.
Our high resolution imaging runs resulted in the detection of only two nearby companions, one in the TOI-4462 system and another in the TOI-5240 system (see Figure 1). The companion to TOI-5240 was detected only by PHARO and is 2.4 away at a position angle of 156 degrees. It is 6.67 magnitudes dimmer in the Br filter, contributing only 0.1% of the total flux of the unresolved system. Even if this companion is a perfectly edge-on, equal mass eclipsing binary, the eclipse depths would be an order of magnitude smaller than the observed transit depths. Since it contributes a negligible amount of light to the overall flux of the system, we chose to neglect this companion in our analysis (eg. Mugrauer & Michel 2020, 2021).
The companion to TOI-4462 A was resolved by both SAI and PHARO at 0.4 separation and a position angle of 225 degrees. It is approximately 2.6 magnitudes dimmer in the and filters, too bright to neglect in our analysis (see §3). However, we remain confident that the Keplerian signals detected in both our photometry and spectroscopy can only be attributed to the brighter, primary star. The transits observed by TESS KeplerCam, and LCO-TEID show no evidence of chromaticity, and the spectral line profiles show no evidence of a secondary set of spectral lines that would produce an apparent RV shift. Hence we are confident that the signal we detect is due to an unresolved transiting companion around the brighter primary star TOI-4462 A.
3 Analysis
We analyzed each system using EXOFASTv2 444https://github.com/jdeast/EXOFASTv2 (Eastman et al. 2019), a publicly available exoplanet fitting suite. EXOFASTv2 is a Differential Evolution Markov Chain Monte Carlo (MCMC) code which globally fits both the star and the companion simultaneously, ensuring a self-consistent set of parameters for the entire system. In each fit, we generate a Spectral Energy Distribution (SED) model for the host star using MESA Isochrones and Stellar Tracks (MIST; Paxton et al. 2011, 2013) in order to fit the host star, while the companion is fit with a standard Keplerian model. Our SED model is fit to broadband archival photometry which we collected from Gaia Data Release 3 (DR3; Gaia Collaboration et al. 2023), 2MASS (Cutri et al. 2003; Skrutskie et al. 2006), and WISE (Wright et al. 2010; Cutri et al. 2012). The Keplerian model for the companion is fit to the TESS and ground based transits as well as the RV data from TRES. For a more detailed explanation of the modeling process, see Eastman et al. (2019).
The fit was generally set up in the same way for each system, except for TOI-4462, which required special consideration due to the presence of a bright nearby companion which we discuss in §3.1. For the other 11 systems, we first compiled the archival photometry for each target in the Gaia , , , 2MASS , , , and WISE , , and bands to construct the SED. We then placed a set of priors on each system based on previous observations, the first of which was a Gaussian prior on the parallax from Gaia DR3 with the Lindegren et al. (2021) correction applied. The parallax uncertainty was added in quadrature with 0.01 to account for any remaining systematic residuals. We also placed a Gaussian prior on the host star metallicity centered on the average value of the TRES-derived metallicity with a prior width of twice the standard deviation. Additionally, we place an upper limit on the V-band extinction along the line-of-site using the dust maps from Schlegel et al. (1998) and Schlafly & Finkbeiner (2011).
In addition to the priors described above, we also fit for a dilution term in each system to account for unresolved contaminants.To do this, we placed a prior of 0% 10% of the contamination ratio reported by the TESS Input Catalog (TIC; Stassun et al. 2018, 2019). While the QLP and SPOC light curves are both already corrected for known contaminants, we still chose to fit for a dilution term as a conservative assumption that the correction applied had a precision of at most 10%. We did this because the contamination ratio reported by the TIC is only an estimate that does not account for the actual point spread functions, as the CCD location and camera were unknown until after the launch of TESS. We also provided each fit with starting points on several parameters from the TIC. Specifically, we adopted the TIC-derived values for the host star’s mass, radius, and effective temperature as well as the companion’s orbital period, time of conjunction, and radius. We retrieved these values from the TESS mission catalog on ExoFOP555https://exofop.ipac.caltech.edu/tess/. We performed a preliminary fit with EXOFASTv2 on each system which included fitting a linear term to the radial velocities in order to account for a long-term drift due to unseen outer companions. In every case except for TOI-4737, this resulted in a slope consistent with zero within 1-sigma, and we subsequently fixed the slope to zero in all subsequent fits for these systems. For TOI-4737, we continued to fit for this long-term trend, and in the final iteration of these fits which we publish here, we found a slope of m s-1 day-1. Each system’s final fit was run to the adopted convergence criteria suggested by Eastman et al. (2019) of at least 1000 independent draws and a Gelman-Rubin statistic . See Table 6 for the priors used, and the median values determined from our analysis.
Four of our fits resulted in bimodal posterior distributions. This typically arises when EXOFASTv2 is unable to distinguish between a host star that is at the end of the main sequence versus the subgiant branch resulting in a high and low stellar mass solutions. Indeed, this was the case in all four bimodal systems presented here (TOI-3577, TOI-4462, TOI-4759, and TOI-5882). We characterized each solution independently by splitting the posterior distributions at the local minimum between the two solutions. We present both solutions for the sake of transparency, but in each case we adopt the higher probability solution as the preferred parameter set. See Table 7 for the priors used, and the median values determined for both solutions of these bimodal systems. Plots of the transit photometry, radial velocities, SED, and MIST evolutionary tracks for each system presented in this work are presented in Figures 6-16.
TOI-2844 | TOI-3122 | TOI-3755 | TOI-4635 | ||
Priors: | |||||
Gaia Parallax (mas) | [1.4759, 0.01696] | [1.962, 0.01778] | [3.0924, 0.01443] | [13.33, 0.02582] | |
Metallicity (dex) | [0.025, 0.198] | [0.3193, 0.1188] | [0.3156, 0.1007] | [-0.1783, 0.1608] | |
V-band extinction (mag) | [0, 0.2725] | [0, 0.4675] | [0, 0.9867] | [0, 0.3959] | |
Dilution in TESS | [0, 0.008335] | [0, 0.027729] | [0, 0.012054] | [0, 0.002608] | |
Primary Star Parameters: | |||||
Mass () | |||||
Radius () | |||||
Luminosity () | |||||
Density (cgs) | |||||
Surface gravity (cgs) | |||||
Effective temperature (K) | |||||
Metallicity (dex) | |||||
Initial metallicity | |||||
Age | Age (Gyr) | ||||
EEP | Equal evolutionary phase | ||||
V-band extinction (mag) | |||||
Distance (pc) | |||||
Companion Parameters: | |||||
Period (days) | |||||
Radius () | |||||
Mass () | |||||
Time of conjunction () | |||||
Optimal conjunction time () | |||||
Semi-major axis (AU) | |||||
Inclination (Degrees) | |||||
Eccentricity | |||||
Argument of periastron (Degrees) | |||||
Equilibrium temperature (K) | |||||
Tidal circularization timescale (Gyr) | |||||
RV semi-amplitude (m/s) | |||||
Radius of planet in stellar radii | |||||
Semi-major axis in stellar radii | |||||
Depth | TESS flux decrement at mid-transit | ||||
Ingress/egress transit duration (days) | |||||
Total transit duration (days) | |||||
Transit impact parameter | |||||
Density (cgs) | |||||
Surface gravity | |||||
Safronov number | |||||
Time of eclipse () | |||||
Total eclipse duration (days) | |||||
Mass ratio | |||||
Separation at mid-transit |
NOTES:
The priors listed at the top of the table are labeled as [mean, standard deviation] if they are Gaussian priors and [lower limit, upper limit] if they are uniform priors.
TOI-4737 | TOI-5240 | TOI-5467 | ||
Priors: | ||||
Gaia Parallax (mas) | [1.7399, 0.01656] | [1.0355, 0.01516] | [1.7819, 0.01704] | |
Metallicity (dex) | [0.3267, 0.1499] | [-0.1571, 0.2089] | [0.3257, 0.1645] | |
V-band extinction (mag) | [0, 1.8386] | [0, 0.4879] | [0, 1.6687] | |
Dilution in TESS | [0, 0.005883] | [0, 0.024878] | [0, 0.007941] | |
Primary Star Parameters: | ||||
Mass () | ||||
Radius () | ||||
Luminosity () | ||||
Density (cgs) | ||||
Surface gravity (cgs) | ||||
Effective temperature (K) | ||||
Metallicity (dex) | ||||
Initial metallicity | ||||
Age | Age (Gyr) | |||
EEP | Equal evolutionary phase | |||
V-band extinction (mag) | ||||
Distance (pc) | ||||
Companion Parameters: | ||||
Period (days) | ||||
Radius () | ||||
Mass () | ||||
Time of conjunction () | ||||
Optimal conjunction time () | ||||
Semi-major axis (AU) | ||||
Inclination (Degrees) | ||||
Eccentricity | ||||
Argument of periastron (Degrees) | ||||
Equilibrium temperature (K) | ||||
Tidal circularization timescale (Gyr) | ||||
RV semi-amplitude (m/s) | ||||
Radius of planet in stellar radii | ||||
Semi-major axis in stellar radii | ||||
Depth | TESSflux decrement at mid-transit | |||
Ingress/egress transit duration (days) | ||||
Total transit duration (days) | ||||
Transit impact parameter | ||||
Density (cgs) | ||||
Surface gravity | ||||
Safronov number | ||||
Time of eclipse () | ||||
Total eclipse duration (days) | ||||
Mass ratio | ||||
Separation at mid-transit |
TOI-3577 | TOI-4462 | ||||
Low-mass solution | High-mass solution | Low-mass solution | High-mass solution | ||
(63.8% probability) | (36.2% probability) | (90.7% probability) | (9.3% probability) | ||
Priors: | |||||
Gaia Parallax (mas) | [2.36440, 0.01487] | [2.53984, 0.03785] | |||
Metallicity (dex) | [-0.0486, 0.0889] | [0.0873, 0.1781] | |||
V-band extinction (mag) | [0, 2.2165] | [0, 0.0949] | |||
Dilution in TESS | [0, 0.032559] | [0, 0.057591] | |||
Primary Star Parameters: | |||||
Mass () | |||||
Radius () | |||||
Luminosity () | |||||
Density (cgs) | |||||
Surface gravity (cgs) | |||||
Effective temperature (K) | |||||
Metallicity (dex) | |||||
Initial metallicity | |||||
Age | Age (Gyr) | ||||
EEP | Equal evolutionary phase | ||||
V-band extinction (mag) | |||||
Distance (pc) | |||||
Companion Parameters: | |||||
Period (days) | |||||
Radius () | |||||
Mass () | |||||
Time of conjunction () | |||||
Optimal conjunction time () | |||||
Semi-major axis (AU) | |||||
Inclination (Degrees) | |||||
Eccentricity | |||||
Argument of periastron (Degrees) | |||||
Equilibrium temperature (K) | |||||
Tidal circularization timescale (Gyr) | |||||
RV semi-amplitude (m/s) | |||||
Radius of planet in stellar radii | |||||
Semi-major axis in stellar radii | |||||
Depth | TESSflux decrement at mid-transit | ||||
Ingress/egress transit duration (days) | |||||
Total transit duration (days) | |||||
Transit impact parameter | |||||
Density (cgs) | |||||
Surface gravity | |||||
Safronov number | |||||
Time of eclipse () | |||||
Total eclipse duration (days) | |||||
Mass ratio | |||||
Separation at mid-transit |
TOI-4759 | TOI-5882 | ||||
Low-mass solution | High-mass solution | Low-mass solution | High-mass solution | ||
(68.2% probability) | (31.8% probability) | (43.5% probability) | (56.5% probability) | ||
Priors: | |||||
Gaia Parallax (mas) | [1.34392, 0.01465] | [2.42207, 0.01753] | |||
Metallicity (dex) | [0.2723, 0.4141] | [0.1400, 0.1900] | |||
V-band extinction (mag) | [0, 0.1507] | [0, 0.8438] | |||
Dilution in TESS | [0, 0.038128] | [0, 0.008565] | |||
Host Star Parameters: | |||||
Mass () | |||||
Radius () | |||||
Luminosity () | |||||
Density (cgs) | |||||
Surface gravity (cgs) | |||||
Effective temperature (K) | |||||
Metallicity (dex) | |||||
Initial metallicity | |||||
Age | Age (Gyr) | ||||
EEP | Equal evolutionary phase | ||||
V-band extinction (mag) | |||||
Distance (pc) | |||||
Companion Parameters: | |||||
Period (days) | |||||
Radius () | |||||
Mass () | |||||
Time of conjunction () | |||||
Optimal conjunction time () | |||||
Semi-major axis (AU) | |||||
Inclination (Degrees) | |||||
Eccentricity | |||||
Argument of periastron (Degrees) | |||||
Equilibrium temperature (K) | |||||
Tidal circularization timescale (Gyr) | |||||
RV semi-amplitude (m/s) | |||||
Radius of planet in stellar radii | |||||
Semi-major axis in stellar radii | |||||
Depth | TESSflux decrement at mid-transit | ||||
Ingress/egress transit duration (days) | |||||
Total transit duration (days) | |||||
Transit impact parameter | |||||
Density (cgs) | |||||
Surface gravity | |||||
Safronov number | |||||
Time of eclipse () | |||||
Total eclipse duration (days) | |||||
Mass ratio | |||||
Separation at mid-transit |
3.1 Multi-star Fitting in EXOFASTv2
As discussed in §2.4, a stellar companion to TOI-4462 was detected 0.4 away, which was blended in all catalog photometry and a significant factor in the dilution of the transit light curves. Given that the probability of a chance alignment is low, and the high Gaia Re-normalized Unit Weight Error (RUWE) of 3.13, we assumed this companion is bound to the primary star. We undid the deblending that SPOC applies to the TESS lightcurves so that we could more accurately model it based on our multi-component SED model. We modeled both stars simultaneously, each with their own MIST evolutionary model, while assuming that the age, initial metallicity, distance, and extinction is the same for both stars. In addition, we modeled a spectral energy distribution for each star, constraining the sum of both stars with the catalog photometry of the unresolved TOI-4462 system, and the difference between the two stars with the AO photometry from PHARO, as shown in Figure 1 (Left). The dilution of the transit light curve was accounted for by fitting for dilution terms that were then constrained by the multi-component SED model, integrated at the transit-observed bands assuming a 5% floor in the theoretical dilution from the model atmospheres. That is, we applied an adaptive prior penalty of
where is the modeled dilution at the current MCMC step and is the SED-derived dilution. This naturally propagates the uncertainty in the stellar properties, accounting for systematics in the theoretical atmospheres, to the light curve de-blending and transit depth.
4 Discussion
The 5 BD-mass companions presented here increase the population of transiting brown dwarfs to over 50. While this number is expected to continue growing, it is worth analyzing the sizable population that has been put together thus far in the context of planet-like and star-like formation. We may find that early trends or features observed in the growing population from previous efforts have been reinforced, or lost significance (perhaps even disappearing) in the wake of new discoveries. One such feature of particular interest is the so-called ”brown dwarf desert” and its potential role in dividing the brown dwarfs into distinct planet-like and star-like groups. We discuss these trends below in §4.1 and §4.2. We also note that one of the BD-hosting stars presented here, TOI-5882, has a significant absorption feature at 6708 Å which we attribute to lithium. We discuss in §4.3 the implications of this, as well as how it affects our determination of the system’s age in.
4.1 The Transiting Brown Dwarf Desert
Perhaps the most discussed feature to emerge from the growing transiting BD population is that of the brown dwarf desert. The phrase was originally coined in the earliest days of exoplanet discovery (Marcy et al. 1997; Latham et al. 1998) in reference to the lack of BD-mass companions discovered by RV surveys at that time. It has since evolved over time with studies like Grether & Lineweaver (2006) which analyzed RV-detected companions with periods less than 5 years and found the ”driest” part of the desert was between . Then Ma & Ge (2014) examined trends in the RV discovered population and refined the measurement to be between and with days.
We now have a more substantial population of transiting BD systems which have precisely measured radii and masses. It’s worth exploring how well the trends found in the the RV discovered sample hold up in the transiting BD regime. Of course, because of the transit probability decreases with period, transiting systems tend to have much shorter orbital periods than their counterparts discovered through direct imaging and RV campaigns. As a result we are largely investigating a different, more limited parameter space than the previous studies of Ma & Ge (2014) for example. We found that the most sparsely populated area of the transiting BD desert appears to be the entire low-mass BD regime ( 42), and the work presented here contributes 1 new BD (TOI-5882) to this underpopulated region (see Figure 2). We also note that the apparent drop off in systems with companions above the substellar limit (80 ) is likely unphysical and is more plausibly due to selection bias since most of these systems have been discovered via exoplanet discovery pipelines. In order to more accurately describe trends emerging near the hydrogen fusion boundary, a more unbiased sample will need to be produced.
Another key trend that was first noted by the Ma & Ge (2014) RV study is in the eccentricity vs distribution. They found that eccentricity decreases as companion mass increases up until 42 , right in the middle of the driest part of the brown dwarf desert. On the other hand, companions more massive than 42 cover a much larger range in eccentricity and little to no correlation with the companion mass. The authors attributed this trend as evidence of a 42 transition point between the planet and star formation mechanisms. As the transiting BD population has developed, several studies have drawn comparisons to these results, some finding evidence for the same trends (Grieves et al. 2021; Henderson et al. 2024a), while others have noted low-mass BDs with higher than expected eccentricities (Page et al. 2024). However, such claims have historically been subject to the small sample size and selection effects that accompany the transiting BD population. Now that this population exceeds 50 systems, we can at least start to alleviate the risks of small number statistics. Figure 3 (Left) shows the eccentricity vs. companion mass distribution for the transiting brown dwarfs, and it is clear that there are more systems above 42 with high eccentricity than there are below. About 30% of systems below 42 have eccentricities compared to about 45% for their higher-mass counterparts. However, we argue that this trend alone is not necessarily supportive of a 42 transition between the planet and stellar formation mechanisms. If the low-mass transiting BDs are indeed dominated by the planet formation mechanism, then they should be subject to the same evolutionary pathways as the hot Jupiters. The hot Jupiter eccentricity distribution has been shown to be most consistent with high eccentricity migration mechanisms and thus are ultimately sculpted by tidal recircularization (Rodriguez et al. 2023; Schulte et al. 2024). We know that this process depends more fundamentally on the mass ratio of the system rather than just the companion mass as evidenced by the tidal recircularization time scale (Equation 2 of Adams & Laughlin 2006). So, if the trend in eccentricity vs companion mass were indeed indicative of a separation between planet-like and star-like formation processes, then we should expect to see the same trends emerge in eccentricity vs mass ratio. Namely, low mass ratio systems should exhibit a much smaller range of eccentricities than their high mass ratio counterparts. However, as shown in Figure 3 (Right), we see the opposite. The eccentricity dichotomy between low and high mass companions seems to disappear when plotted against mass ratio. We therefore argue that the eccentricity vs companion mass distribution of transiting companions does not support a 42 transition point.
4.2 Transiting Brown Dwarf Metallicities
Eccentricity is likely not the only parameter that could offer insight into at which critical companion mass the dominant formation mechanism turns over from planet-like to star-like. For nearly two decades we have known of the giant planet-metallicity correlation, in which hot Jupiter hosting stars tend to be more metal rich than their counterparts with no discovered planets (Gonzalez 1997; Santos et al. 2003; Fischer & Valenti 2005). If the low-mass transiting BDs are predominantly forming in the same way as the hot Jupiters, then we should expect their host stars to exhibit the same metallicity enhancement when compared to the high-mass BD hosts. Schlaufman (2018) tested this hypothesis using the metallicities of transiting companions in the range to show that transition between core accretion and fragmentation may be as low as . However, at the time of this study, there were only 27 transiting companions known between limiting the ability to probe potential higher-mass transition points. Now that we have access to significantly more systems in this mass regime, we can better probe the same 42 transition. In Figure 4 we show a preliminary look at testing this hypothesis. Qualitatively, it appears that the lower-mass companions ( ) preferentially orbit more metal-rich host stars. However, a two-sample Kolmogovorov-Smirnov test yields a -value of 0.35, too high to reject the null hypothesis that the high mass and low mass samples are drawn from the same underlying distribution.
We note also a few important caveats for interpreting the metallicity distribution. First, we chose a 42 cutoff for historical reasons in order to compare to the original hypothesis presented by Ma & Ge (2014) as well as the eccentricity distribution presented in §4.1. It may be that a more appropriate boundary separating the population will be found after a more comprehensive analysis which may be in better agreement with the lower-mass Schlaufman (2018) transition. With just 26 companions below 42 in our sample, we chose note to investigate possible lower-mass transitions. A more in-depth analysis including the population of giant planets will need to be done to more precisely probe a lower mass transition. An unbiased sample of the population of companions across the substellar limit will also be required to better combat the selection effects currently affecting this population. We note also that the metallicities presented here are the reported values from each system’s discovery which have been measured using a variety of different techniques and therefore the underlying biases affecting each measurement are not explored here.
4.3 Lithium Detected in TOI-5882
During our analysis of TOI-5882’s spectra, we found a significant absorption feature at 6708 Å which we attribute to the lithium doublet. We measured the equivalent width of this feature using the specutils (Earl et al. 2022) package in Python (see Figure 5). To perform this measurement, we first co-added all of the observed TRES spectra, after correcting each for the RV shift, to increase the signal-to-noise resolution (SNR) of the Li feature. The resulting co-added echelle order containing Li has an SNR of 56.8. Then, we defined a 1.4 Å region centered on the rest wavelength of the Li doublet at 6707.844 Å to measure the equivalent width. The resulting equivalent width is mÅ, where the uncertainty was estimated using Equation 6 in Cayrel (1988).
The presence of lithium in stars is typically interpreted as an indicator of youth. This is due to the temperatures and pressures in the core being sufficiently high to destroy Li, which results in Li visible on the stellar surface slowly depleting as transport occurs between the core and the surface of the star (Soderblom et al. 2014). Despite this, we claim that TOI-5882 is likely not a young star, since we found no other signs of youth. Specifically, we found no evidence of rapid rotation, comoving companions, or an infrared excess. To verify this, we performed a period search on each sector of TOI-5882’s TESS light curves to characterize the rotation of the host star. Since young stars are typically born rapidly rotating and gradually spin down over time, an age can often be inferred from a star’s rotation period if it is below the Kraft break (Bouma et al. 2023). We found a significant peak in the periodogram at 9.6 days, however, we are hesitant to adopt this as the true rotation period since periodicity beyond 1/3 of a TESS observing sector ( 9 days) can be unreliable due aliases induced by the TESS observing strategy and processing of light curves. Nevertheless, this rotation period would still be much too long to be consistent with a young age given this star’s effective temperature (Bouma et al. 2023).
We also searched for nearby comoving stars using FriendFinder 666https://github.com/adamkraus/comove (Tofflemire et al. 2021) since their presence would indicate that TOI-5882 and its hypothetical nearby comovers have not yet dispersed from their birth location and hence would be young. FriendFinder identifies all nearby sources that fall within a selected search radius, and calculates the predicted tangential velocity for each source, assuming that they have Galactic velocity components (,,) identical to TOI-5882. This predicted is then compared to the true which is derived from Gaia proper motions. Using a physical search radius of 30 pc around TOI-5882 and a difference between the predicted and measured of km s-1 we find no evidence that TOI-5882 is part of a comoving group.
Finally, we looked for an infrared excess since young stellar objects that retain a circumstellar disk show increased emission at infrared wavelengths (Cotten & Song 2016). We rule out an infrared excess for TOI-5882 via our SED fitting in the global analysis where we see no significant infrared emission above the blackbody model in any of the WISE , , and bandpasses. While the lack of these additional youth indicators do not completely rule out the possibility of a young host star, we believe it is more likely that TOI-5882 is a late main-sequence star, as indicated by our most probable EXOFASTv2 solution. This older age could then imply that the presence of Li is due to the infall of planetary material onto the host star. For a deeper dive into the origin of Li in TOI-5882, including its potential as system that has undergone a planetary engulfment, we alert the reader to Kotten et al. (in prep).
5 Conclusions
In this paper, we present the discovery of 11 new transiting companions from the TESS mission. We collected photometric time-series, spectroscopic, and high resolution imaging follow-up as a part of the TESS Follow-up Observing Program (TFOP) to rule our false positives and further characterize each system. Using EXOFASTv2, we performed a global fit on each system using the space and ground-based transits, spectroscopic RVs, and archival photometry to characterize both the host stars, and their transiting companions. We found that 5 of these systems are brown dwarfs () and 6 of them are very low mass stars from . This contribution to the transiting brown dwarf population increases it to 54 systems, a milestone that represents the population outgrowing burdens of small sample statistics.
Using this population that TESS has rapidly developed to a significant size, we offered some initial insight into the features that have started to appear. We revisit the idea of the ”brown dwarf” desert for the short orbital periods probed by the transit method. We revisited the eccentricity-mass distribution that has been claimed as evidence of a 42 transition between planet and star formation and showed that this trend does not seem to hold in eccentricity versus mass-ratio, calling into question whether eccentricity truly does offer insight into the formation mechanisms behind these rare objects. We also examined the metallicity distribution of transiting BD host stars for the first time, and find that a 42 transition does not divide the population into two distinct populations with any statistical significance.
Finally, we noted the presence of Li in the spectrum of TOI-5882, the host star of our lowest mass BD. We measured the equivalent width of the Li line and search for other signs of youth. Seeing no evidence of rapid rotation, nearby comoving stars, or infrared excess, we adopt the age provided by our global EXOFASTv2 fit and do not interpret the presence of Li as a sign of youth. Instead we noted that the Li may actually be a signature of engulfed planetary material, and that more work will be required to explore this hypothesis.
Acknowledgements
This paper includes data collected by the TESS mission that are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This research has made use of the NASA Exoplanet Archive and the Exoplanet Follow-up Observation Program (ExoFOP; DOI: 10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. We acknowledge financial support from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 and the ERDF “A way of making Europe” through project PID2021-125627OB-C32, and from the Centre of Excellence “Severo Ochoa” award to the Instituto de Astrofisica de Canarias. This work makes use of observations from the LCOGT network. Part of the LCOGT telescope time was granted by NOIRLab through the Mid-Scale Innovations Program (MSIP). MSIP is funded by NSF. This paper is based on observations made with the Las Cumbres Observatory’s education network telescopes that were upgraded through generous support from the Gordon and Betty Moore Foundation. Based on observations obtained at the Hale Telescope, Palomar Observatory, as part of a collaborative agreement between the Caltech Optical Observatories and the Jet Propulsion Laboratory operated by Caltech for NASA. DRC and CAC acknowledge partial support from NASA Grant 18-2XRP18_2-0007. Based in part on observations obtained at the Southern Astrophysical Research (SOAR) telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the U.S. National Science Foundation NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This work is partly supported by JSPS KAKENHI Grant Number JP24H00017, JP24K00689 and JSPS Bilateral Program Number JPJSBP120249910. This paper is based on observations made with the MuSCAT2 instrument, developed by ABC, at Telescopio Carlos Sánchez operated on the island of Tenerife by the IAC in the Spanish Observatorio del Teide.
NV is supported by the NASA FINESST program. The postdoctoral fellowship of KB is funded by F.R.S.-FNRS grant T.0109.20 and by the Francqui Foundation. KAC and CNW acknowledge support from the TESS mission via subaward s3449 from MIT. I.A.S. acknowledges the support of M.V. Lomonosov Moscow State University Program of Development. F. M. acknowledge support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) through grant PID2023-152906NA-I00. LM acknowledges financial contribution from PRIN MUR 2022 project 2022J4H55R.